OpenStudy (anonymous):

what is the function rule that represents the relationship found in ordered pairs

OpenStudy (anonymous):

(-5,22),(10,-23),(4,-5),(-7,28),(6,-11)

OpenStudy (compassionate):

\(\large Since~you~seem~new~here:\\\bf\Huge{~~\color{red}{\boxed{W}}\color{#FF9200}{\boxed{E}}\color{#FFDB00}{\boxed{L}}\color{#B6ff00}{\boxed{C}}\color{#00ff49}{\boxed{O}}\color{#00DBff}{\boxed{M}}\color{#B600ff}{\boxed{E}}\\~~~~~~~~~~~~~~~~~~\boxed{T}\boxed{O}\\~~~~~~~~~~~\color{#0092ff}{\boxed{O}}\color{#0092ff}{\boxed{P}}\color{#0092ff}{\boxed{E}}\color{#0092ff}{\boxed{N}}\color{#7cc517}{\boxed{S}}\color{#7cc517}{\boxed{T}}\color{#7cc517}{\boxed{U}}\color{#7cc517}{\boxed{D}}\color{#7cc517}{\boxed{Y}}\color{#7cc517}{\boxed{!}}} \\ \large\bf You~can~read~the~\underline{\href{ /code-of-conduct }{Code~of~Conduct}}~here. • \\ \it \large Now~let's~try~to~find~an~answer~to~your~problem~together.\)\ Are you aware of how to find a function? It's kind of like a trial and error process. (-5,22),(10,-23),(4,-5),(-7,28),(6,-11) x = -5, 10, 4, -7, 6 y = 22, -23, -5, 28, -11. The next phase is to put them in order (least > greatest). -7, -5, 4, 6, 10 -23, -11, -5, 22, 28 _____________________ Now, can you see any pattern? Comment back and let me know. (As x increases, y increases by ?)

OpenStudy (anonymous):

i cant see a pattern im so lost.

OpenStudy (compassionate):

Take -7 and find the difference between it and -23. Find the difference of -5 and -11. For example, the difference between 10 and 28 is 18.

OpenStudy (anonymous):

-30,-16,-1,-16, and -18

OpenStudy (compassionate):

you might wanna redo that :/

OpenStudy (anonymous):

i subtract right

OpenStudy (compassionate):

When dealing with negaive numbers it's a bit different. You want to just count the numbers between.

OpenStudy (anonymous):

16,6,1,16,18

OpenStudy (compassionate):

Assuming it is linear (4,-5) (6, -11) slope =-6/2 =-3 y+5 =-3(x-4) y+5 =-3x +12 y=-3x +7 it works for all the other points